OpenStudy (anonymous):

Mr. Harrow has eight boys and six girls in his Honors Precalculus class. If he randomly chooses two students, one at a time, what is the probability that they are both girls?

OpenStudy (anonymous):

how many total students?

OpenStudy (anonymous):

8+6 is 14

OpenStudy (anonymous):

no what is the probability that both are girls?

OpenStudy (anonymous):

right. now we are looking for the prob of picking a girl 1st. waht do you think that prob is?

OpenStudy (anonymous):

i know its 15/91 but idk you i am supposed to get that?

OpenStudy (anonymous):

get the prob of picking a girl on the 1st try and I will take you thru it

OpenStudy (anonymous):

1/6?

OpenStudy (anonymous):

6 out of 14: first girl 5 out of 14 :second girl then you multiply them : 6*5\(14*14)= 30\196 simplify, it makes 15\98

OpenStudy (anonymous):

ohh thank you!

OpenStudy (anonymous):

6 girls and 14 students. 6/14=3/7 now on the next oick there are only 5 girls and only 13 students 5/13 multiply these 2 togethher to get 15/91

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

okay i get it now, i was doing 1/6 times 1/5

OpenStudy (anonymous):

great